HUBBLE SPACE TELESCOPE DAILY REPORT #2320 PERIOD COVERED: 0000Z (UTC) 02/02/99 - 0000Z (UTC) 03/03/99 Daily Status Report as of 062/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed STIS/CCD 8080 (STIS Spectroscopic CCD Flatfielding -- Revised) The Space Telescope Imaging Spectrograph (CCD) was used to make CCD flat fielding calibration observations in the spectroscopic mode. The observations were completed as planned, and no anomalies were reported. 1.3 Completed WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxies 3C124.0 and 3C107.0. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena and they may be used as probes of their environments. They are unique probes of the early Universe. The observations were completed as planned, and no problems were reported. 1.4 Completed Five Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.5 Completed WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 6699 (The Globular Cluster System of the Dwarf Elliptical System Galaxies NGC-185 and NGC 205) The WF/PC-2 was used to make observations of the globular cluster system around the nucleus of dwarf galaxy NGC-205. The propose is to carry out a general reconnaissance of the properties of the globular clusters. The observations were completed as planned, and no anomalies were reported. 1.8 Completed Two Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The second iteration of this proposal did not execute due to the acquisition failure described in HSTAR 7024 and 2.1, keeping the take data flag down. Otherwise, the observations were completed as scheduled, and no further problems were reported. 1.9 Completed WF/PC-2 6777 (A Counterpart Search for SGR 0526-66) The WF/PC-2 was to be used to search for the optical counter part of Short Gamma Ray Repeater (SGR) 0526-66 (a.k.a., N49-052601). Soft gamma-ray repeaters are a class of bursting X-ray sources which are characterized by short durations, relatively soft spectra, and lack of spectral evolution during outburst. This proposal did not execute due to the acquisition failure described in HSTAR 7024 and 2.1, keeping the take data flag down. 1.10 Completed WF/PC-2 7405 (A WF/PC-2 Search for Surviving Binary Companions in SN Ia Remnants) The WF/PC-2 was used to search for binary companions in the class Ia super nova remnant SN1006-Cen. Type Ia supernovae {SNe Ia} are widely believed to be explosions of white dwarfs undergoing binary mass accretion. The observations were completed as planned, and no anomalies were reported. 1.11 Completed WF/PC-2 6816 (The Small Scale Structure and Symmetries in Proto-Planetary and Young Planetary Nebulae) The WF/PC-2 was used to make observations of FROSTYLEO in order to study its small scale structure and summetries. Proto-planetary and young planetary nebulae show a dazzling variety of non-spherical morphologies, whereas the circumstellar envelopes of the AGB red giant stars from which they evolve are mostly round. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 6461 (High-Resolution Next Bright Galactic Nova) The WF/PC-2 was used to make high resolution images of the bright galactic nova U-Scorpio after it erupted to search for fluorescence and reflection light echoes from, and small-scale structure in, the nova circumstellar gas and dust ejecta. These observations were a target of opportunity, and a new SMS had to be built. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 8391 (Calibration of the MOC on MGS) The WF/PC-2 was used to make images of the planet Mars. WF/PC-2 images of Mars obtained during the first week of March 1999 will be used to help establish the absolute calibration of the Mars Orbiter Cameras {MOC} on the Mars Global Surveyor {MGS}. The observations were completed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 As described in HSTAR 7024, the acquisition at 061/1015Z failed (not reported yesterday). The proposals described in 1.8 and 1.9 were affected. Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 19 Successful : 19 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared once. Because of the acquisition failure detailed in 2.1, an SI C&DH error flag was cleared (using ROP NS-5) and the STIS flight software error counter was reset (using ROP NS-12). The SSA transmitter was turned on and off once per ROP IC-2. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH